OpenStudy (anonymous):

Simplify the trigonometric expression? sin^2 θ/ (1 - cos θ)

OpenStudy (raden):

sin^2 θ = 1 - cos^2 θ

OpenStudy (raden):

the use the difference square formula : a^2 - b^2 = (a+b)(a-b) so, sin^2 θ/ (1 - cos θ) = (1 - cos^2 θ)/ (1 - cos θ) = (1 + cos θ) (1 - cos θ)/ (1 - cos θ) cancel out the same term, and get ur answer

OpenStudy (raden):

* then

OpenStudy (anonymous):

thank you very much! :)

OpenStudy (raden):

wlcm